OpenStudy (gabrielah96):

prove the identity (1-3cosx-4cos^2x)/sin^2x=(1-4cosx)/1-cosx

OpenStudy (anonymous):

Factor

OpenStudy (anonymous):

\[-(4\cos^2x + 3cosx -1)/(\sin^2x) = (4cosx -1)/(cosx + 1)\]